OpenStudy (anonymous):

Multiply. (a+1/5)(a+1/2)

OpenStudy (amistre64):

sorry caps werent meant to be on

OpenStudy (amistre64):

do you remember how to multiply?

OpenStudy (anonymous):

a^2+7/10a+1/10

OpenStudy (amistre64):

a+1/5 a+1/2 -------- a^2 +a/5 a/2 + 1/10 --------------- a^2 +7a/10 + 1/10

OpenStudy (anonymous):

a=-1/2 a=-1/5 is what I am getting.

OpenStudy (amistre64):

you arent told to solve for anything; just to rewrite it

OpenStudy (anonymous):

you said multiply, not solve

OpenStudy (anonymous):

multiply is all it says

OpenStudy (anonymous):

nope, we multiplied it, like you said solving would be your answer

OpenStudy (anonymous):

a=-1/2 or a=-1/5

OpenStudy (amistre64):

\[\frac{1}{10}(10a^2+7a+1)\]

OpenStudy (anonymous):

so a^2+7a/10+1/10 is the answer thanks